(Common stock valuation​) Abercrombie​ & Fitch's common stock pays a dividend of $ 3.00$3.00. It is...

(Common

stock

valuation​)

Abercrombie​ & Fitch's common stock pays a dividend of

$ 3.00$3.00.

It is currently selling for

$ 37.77

If the​ firm's investors require a return of

13

percent on their investment from buying Abercrombie​ & Fitch​ stock, what growth rate would Abercrombie​ & Fitch have to provide the​ investors?
Answer #1

value = dividend next year/(Required return - growth rate)

=>

37.77 = 3*(1+g)/(0.13-g)

=>

37.77*0.13 - 37.77g = 3+3g

=>

g = 1.9101/40.77

= 4.69%